# MambaFlow: A Mamba-Centric Architecture for End-to-End Optical Flow Estimation

### v1

@@ 3.5 Global Matching
After extracting the enhanced features $\boldsymbol{\mathit{F}}_{q}$ and $\boldsymbol{\mathit{F}}_{v}$ from our FEM module, we construct a 4D cost volume of size $H\times W\times H\times W$ by computing the dot-product similarity between all pixel pairs from both features. The construction process is expressed as:
$\mathbf{C}=\frac{\boldsymbol{\mathit{F}}_{q}\,\boldsymbol{\mathit{F}}_{v}^{\top}}{\sqrt{D}},$ (20)
where $D$ is the feature dimension. The constructed 4D cost volume can be interpreted as a collection of $H\times W$ 2D cost maps, where each map evaluates the feature similarity between a single pixel in $\boldsymbol{\mathit{F}}_{q}$ and all pixels in $\boldsymbol{\mathit{F}}_{v}$. To obtain a matching distribution, a softmax operation is applied across the last two dimensions of the cost volume:
$\mathbf{M}=\text{softmax}(\mathbf{C}).$ (21)
To estimate the optical flow, the correspondence $\hat{\mathbf{G}}$ is obtained by taking a weighted average of the 2D pixel grid coordinates $\mathbf{G}\in\mathbb{R}^{H\times W\times 2}$ in the target image with the matching distribution $M$:
$\hat{\mathbf{G}}=\mathbf{M}\mathbf{G}.$ (22)
Finally, the optical flow $\mathbf{V}_{\text{original}}$ is calculated as the difference between the corresponding pixel coordinates:
$\mathbf{V}_{\text{original}}=\hat{\mathbf{G}}-\mathbf{G}.$ (23)
@@
3.6 Flow Propagation Mamba
To address the issue of occluded regions or pixels outside the image boundaries, we introduced a self-similarity-based module called the flow propagation Mamba (FPM) module (as illustrated in Fig. 4), which leverages feature information from the source image to capture pixel-level similarity relationships. Since adjacent pixels typically exhibit similar features, they demonstrate higher similarity, thus providing a more robust solution for optical flow estimation. Specifically, by inputting $\boldsymbol{\mathit{F}}_{q}$ into the Mamba block, we enhance the features in occluded regions and further extract the weights of these occluded areas. This helps optimize the final optical flow estimation, thereby improving accuracy and robustness in complex scenarios. The formula is as follows:
$\mathbf{V}_{\text{final}}=\text{softmax}\left(\text{Mamba}(\boldsymbol{\mathit{F}}_{q})\right)\mathbf{V}_{\text{initial}},$ (24)
where $\mathbf{V}_{\text{initial}}\in\mathbb{R}^{H\times W\times 2}$ represents the optical flow predicted by the global matching.
Figure 4: The architecture of flow propagation Mamba (FPM) module.
Figure 5: Visual results on Sintel test set.
@@ 3.6 Flow Propagation Mamba
MambaFlow (ours) 1.60 2.78 8.53 8.9M 113.5ms
Table 1: MambaFlow achieves at least 18% faster inference on the KITTI (test) dataset compared to other state-of-the-art methods with competitive performance. All compared methods were evaluated under the same training configuration, and a unified ”C+T+S+K+H” training strategy was adopted to ensure the fairness of the results. For performance measurements, we used the normalized setting of batch size 1 and input resolution 540 × 960. In addition, all inference times are measured on the RTX 3090 graphics card. The significant performance improvement highlights MambaFlow’s advantages in speed and efficiency. It is worth noting that while maintaining competitive accuracy on the Sintel and KITTI datasets, we have successfully achieved a significant improvement in speed, further verifying the superiority of MambaFlow in the speed-accuracy balance.
@@ 4 Experiments
full 6.45 - 2.23 - 3.41 - 8.1 -
w/o self 6.89 $\uparrow$6.82.30 $\uparrow$3.13.49 $\uparrow$2.36.9 $\downarrow$14.8%
w/o cross 11.10 $\uparrow$72.13.78 $\uparrow$69.55.16 $\uparrow$51.34.4 $\downarrow$45.7%
w/o MLP 7.13 $\uparrow$10.52.41 $\uparrow$8.13.66 $\uparrow$7.35.2 $\downarrow$35.8%
w/o pos 6.65 $\uparrow$3.12.29 $\uparrow$2.73.48 $\uparrow$2.18.1 -
Table 2: FEM components. Ablation study on individual components, revealing that Cross-Mamba contributes the most.
layers Sintel (train) KITTI (train) Param (M)
@@
4 1.41 2.79 9.53 27.15 5.1
6 1.37 2.72 9.16 25.38 6.9
8 1.36 2.63 8.54 24.97 8.9
10 1.35 2.68 8.51 24.67 10.8
12 1.36 2.55 8.10 23.91 12.7
24 1.34 2.50 7.96 22.97 23.4
Table 3: Number of Mamba layers. Using 8 layers of Mamba strikes an optimal balance between accuracy and parameter efficiency.
Implementation Details.
Our model is constructed using the PyTorch library and trained on 4 NVIDIA GeForce RTX 3090 GPUs. Following the training strategy from previous works [36, 31], we use the AdamW optimizer. The model is first pre-trained on the Chairs dataset for 100k iterations with a batch size of 16 and a learning rate of $4\times 10^{-4}$. Then, it is fine-tuned on the Things dataset for 200k iterations with a batch size of 8 and a learning rate of $2\times 10^{-4}$. During this process, we select the best-performing model on the Things dataset and fine-tune it for an additional 800k iterations to validate the ablation study. Subsequently, the model fine-tuned for 200k iterations on the Things dataset is used as the initial model and further trained on the Sintel dataset, keeping the same hyperparameters as those used for the Things dataset, including a batch size of 8 and a learning rate of $2\times 10^{-4}$. Finally, we continue training on the KITTI-2015 training set for 200k iterations with the batch size and learning rate unchanged.
@@ 4.1 Comparison with State-of-the-Arts
MambaFlow achieved an EPE of 1.60 on the Sintel Clean dataset and an EPE of 2.78 on the Final dataset. Fig. 5 presents a qualitative analysis of optical flow estimation, comparing MambaFlow with other state-of-the-art approaches on the Sintel test set. The results show that MambaFlow’s performance on Clean and Final is at the same level as current mainstream methods such as RAFT and GMA, demonstrating its strong adaptability in both simple and complex scenarios.
@@ 4.2 Ablation Analysis
Components of FEM.
We ablate different Fearture Enhancement Mamba components in Table 2. Among them, Cross-Mamba contributes the most as it models the mutual relationship between two features, a relationship that is otherwise missing in the features extracted from the convolutional backbone. Self-Mamba aggregates contextual cues within the same feature, resulting in an additional performance improvement, and positional information introduces location dependence into the matching process, helping to mitigate ambiguities present in feature-similarity-based matching alone. Removing the Multi-Layer Perceptron (MLP) reduces the number of parameters, but also results in a slight decrease in performance.
Number of FEM layers.
We conduct an ablation study on the number of FEM layers, as shown in Table 3. The experiment compares the impact of varying the number of FEM layers on overall performance. Among all the tested configurations, the FEMG module with 8 Mamba layers exhibits the best performance.
@@
The experimental results in Table 4 show that with the increase in the number of training iterations, our framework has achieved certain improvements in performance. This shows that more training iterations can effectively improve the performance of the model, making it more stable and accurate in subsequent tasks.
training Things (val) Sintel (train) KITTI (train)
@@
200K 3.04 1.19 2.60 9.02 25.92
400K 2.81 1.14 2.57 8.76 25.18
600K 2.73 1.10 2.52 8.14 24.18
800K 2.61 1.09 2.56 7.57 23.51
Table 4: Training iterations. MambaFlow benefits from more training iterations.
@@ 5 Conclusion
In this paper, we explore for the first time the application of the advanced Mamba in optical flow estimation, and then propose a simple, accurate and efficient Mamba-centric optical flow method, MambaFlow, for end-to-end flow prediction. It consists of two components built on Mamba architecture: the FEM module and FPM module. The FEM is proposed for accurate and efficient feature enhancement, incorporating the Self-Mamba, Cross-Mamba mechanisms to tackle the issue of inadequate feature representation. Furthermore, the FPM that employs feature self-similarity to direct the ultimate flow is crafted for the flow propagation, addressing the challenge of traditional recurrent refined methods. Our method achieves comparable accuracy to state-of-the-art algorithms while significantly reducing inference time, with a runtime of 0.113 seconds—18% faster than GMFlow and 10 times faster than MS-RAFT+. We attribute this to the high accuracy and efficiency of Mamba in capturing features with local correlation while maintaining its global information. As far as we know, this is the first Mamba-centric architecture designed for end-to-end optical flow estimation. We are assured that additional variants of the Mamba-centric methods may emerge , possessing significant potential to boost both precision and velocity in the future.

### v2

@@ 3.5 Global Matching
After extracting the enhanced features $\boldsymbol{\mathit{F}}_{q}$ and $\boldsymbol{\mathit{F}}_{v}$ from our PolyMamba module, we construct a 4D cost volume of size $H\times W\times H\times W$ by computing the dot-product similarity between all pixel pairs from both features. The constructed 4D cost volume can be interpreted as a collection of $H\times W$ 2D cost maps, where each map evaluates the feature similarity between a single pixel in $\boldsymbol{\mathit{F}}_{q}$ and all pixels in $\boldsymbol{\mathit{F}}_{v}$. To obtain a matching distribution, a softmax operation is applied across the last two dimensions of the cost volume. The construction process is expressed as:
$\mathbf{M}=\text{softmax}(\mathbf{\frac{\boldsymbol{\mathit{F}}_{\mathit{q}}\,\boldsymbol{\mathit{F}}_{\mathit{v}}^{\top}}{\sqrt{D}}}).$ (16)
where $D$ is the feature dimension. To estimate the optical flow, the correspondence is obtained by taking a weighted average of the 2D pixel grid coordinates $\mathbf{G}\in\mathbb{R}^{H\times W\times 2}$ in the target image with the matching distribution $M$. Finally, the optical flow $\mathbf{V}_{\text{initial}}$ is calculated as the difference between the corresponding pixel coordinates:
$\mathbf{V}_{\text{initial}}=\mathbf{M}\mathbf{G}-\mathbf{G}.$ (17)
@@
3.6 PulseMamba
To effectively address the challenges of occlusions and boundary conditions in optical flow estimation, we propose a novel module named PulseMamba. PulseMamba integrates cost volume information and motion features, achieving dynamic and context-aware optical flow propagation through iterative refinement.
Given the initial optical flow $\mathbf{V}_{\text{initial}}$, the source image features $\boldsymbol{F}_{q}$, and the 4D cost volume (denoted as $\mathbf{C}$), PulseMamba proceeds as follows:
First, motion features are extracted by encoding the initial flow and the 4D cost volume:
$\boldsymbol{M}=\text{MotionEncoder}(\mathbf{V}_{\text{initial}},\mathbf{C}),$ (18)
PulseMamba employs multiple stacked Mamba layers for iterative feature refinement. At each iteration $i$, the input is formed by concatenating the image features, motion features, and current flow prediction, and the flow is updated as:
$\mathbf{V}^{(i)}=\text{Mamba}^{(i)}\left([\boldsymbol{F}_{q};\;\boldsymbol{M};\;\mathbf{V}^{(i-1)}]\right),$ (19)
where $\mathbf{V}^{(0)}=\mathbf{V}_{\text{initial}}$.
The output flow from each iteration becomes the input for subsequent refinement. After $N$ iterations, the final refined optical flow is obtained:
$\mathbf{V}_{\text{final}}=\mathbf{V}^{(N)}.$ (20)
Through this iterative and context-aware refinement, PulseMamba significantly improves the accuracy and robustness of optical flow estimation, particularly under complex conditions involving occluded regions and ambiguous boundaries.
Table 1: Quantitative results on Sintel and KITTI. MambaFlow achieves at least 18% faster inference on the KITTI (test) dataset compared to other state-of-the-art methods with competitive performance. All compared methods were evaluated under the same training configuration, and a unified "C+T+S+K+H" training strategy was adopted to ensure the fairness of the results. For performance measurements, we used the normalized setting of batch size 1 and input resolution 540 × 960. In addition, all inference times are measured on the RTX 3090. The significant performance improvement highlights MambaFlow’s advantages in speed and efficiency. It is worth noting that while maintaining competitive accuracy on the Sintel and KITTI datasets, we have successfully achieved a significant improvement in speed, further verifying the superiority of MambaFlow in the speed-accuracy balance.
@@ 3.6 PulseMamba
MambaFlow (ours) 1.43 2.45 8.06 11.2M 113.5ms
@@ 4 Experiments
Implementation Details.
Our model is constructed using the PyTorch library and trained on 8 NVIDIA GeForce RTX 3090 GPUs. Following the training strategy from previous works [54, 45], we use the AdamW optimizer. The model is first pre-trained on the Chairs dataset for 100k iterations with a batch size of 16 and a learning rate of $4\times 10^{-4}$. Then, it is fine-tuned on the Things dataset for 200k iterations with a batch size of 8 and a learning rate of $2\times 10^{-4}$. During this process, we select the best-performing model on the Things dataset and fine-tune it for an additional 800k iterations to validate the ablation study. Subsequently, the model fine-tuned for 200k iterations on the Things dataset is used as the initial model and further trained on the Sintel dataset, keeping the same hyperparameters as those used for the Things dataset, including a batch size of 8 and a learning rate of $2\times 10^{-4}$. Finally, we continue training on the KITTI-2015 training set for 200k iterations with the batch size and learning rate unchanged.
Figure 3: Visual results on Sintel test set.
Table 2: PolyMamba components. Ablation study on individual components, revealing that Cross-Mamba contributes the most.
@@
full 2.91 - 1.23 - 2.76 - 11.2 -
w/o self 3.11 $\uparrow$6.91.27 $\uparrow$3.32.82 $\uparrow$2.28.9 $\downarrow$24.6%
w/o cross 5.01 $\uparrow$72.22.08 $\uparrow$69.14.18 $\uparrow$51.45.4 $\downarrow$54.2%
w/o MLP 3.22 $\uparrow$10.71.33 $\uparrow$8.12.96 $\uparrow$7.28.2 $\downarrow$30.5%
w/o pos 3.00 $\uparrow$3.11.26 $\uparrow$2.42.82 $\uparrow$2.211.2 -
Table 3: Number of PolyMamba blocks. Using 8 blocks of Mamba strikes an optimal balance between accuracy and parameter efficiency.
blocks Sintel (train) KITTI (train) Param (M)
@@
4 1.28 2.94 10.03 29.18 7.2
6 1.24 2.87 9.64 27.23 9.2
8 1.23 2.77 8.99 26.84 11.2
10 1.22 2.82 8.96 26.52 13.2
12 1.23 2.69 8.53 25.70 15.2
24 1.21 2.63 8.38 24.69 27.3
@@ 4.1 Comparison with State-of-the-Arts
MambaFlow achieved an EPE of 1.43 on the Sintel Clean dataset and an EPE of 2.45 on the Final dataset. Fig. 3 presents a qualitative analysis of optical flow estimation, comparing MambaFlow with other state-of-the-art approaches on the Sintel test set. The results show that MambaFlow’s performance on Clean and Final is at the same level as current mainstream methods such as RAFT and GMA, demonstrating its strong adaptability in both simple and complex scenarios.
@@ 4.2 Ablation Analysis
Components of PolyMamba.
We ablate different PolyMamba components in Table 2. Among them, Cross-Mamba contributes the most as it models the mutual relationship between two features, a relationship that is otherwise missing in the features extracted from the convolutional backbone. Self-Mamba aggregates contextual cues within the same feature, resulting in an additional performance improvement, and positional information introduces location dependence into the matching process, helping to mitigate ambiguities present in feature-similarity-based matching alone. Removing the MLP reduces the number of parameters, but also results in a slight decrease in performance.
Number of PolyMamba blocks.
We conduct an ablation study on the number of PolyMamba blocks, as shown in Table 3. The experiment compares the impact of varying the number of PolyMamba blocks on overall performance. Among all the tested configurations, the PolyMamba with 8 blocks achieves a good balance between accuracy and parameters.
Number of PulseMamba iterations.
As shown in Table 5, we investigated the impact of different iterations of PulseMamba on EPE on the Sintel dataset. As the number of iterations increases from 1 to 3, the overall EPE gradually decreases from 0.77 to 0.65; the error of the matched area decreases from 0.58 to 0.49, and the error of the unmatched area decreases significantly from 3.68 to 3.16, with the most obvious improvement. This trend shows that more iterations can improve the prediction accuracy of the occluded area.
Weight sharing vs. non-sharing comparison.
We compared Mamba modules with shared and independent weights at the same number of iterations. As shown in Table 5, the non-shared scheme achieved slightly better results on all metrics, matching and non-matching metrics. This is because the parameters of each layer can more flexibly adapt to feature changes between iterations. For example, in two iterations, shared weights reduced the model size by 7.5% (from 12.0 to 11.1), but caused the error to increase by 5.9% (from 0.68 to 0.72). These findings suggest that for complex matching tasks, it is reasonable to accept a moderate increase in the number of parameters due to the improvement in feature representation ability.
@@
As shown in Table 5, Mambaflow has achieved certain improvements in performance as the number of training iterations increases. This shows that more training iterations can effectively improve the performance of the model, making it more stable and accurate in subsequent tasks.
Table 4: Ablations on PulseMamba designs.
PulseMamba iterations share Mamba? Sintel (train, clean)
all matched unmatched
0 2.97 0.68 15.75
1 0.77 0.58 3.68
@@
[…]
Table 5: Training iterations.
[…]
@@ 5 Conclusion
As Mamba continues to gain significant progresses in the field of computer vision, this paper explores for the first time to leverage it to estimate optical flow in an end-to-end paradigm. We suggest a Mamba-centric optical flow method, MambaFlow, which contributes two core Mamba-based components: PolyMamba and PulseMamba, proposed to accurately and efficiently enhance feature representation and direct the ultimate flow, respectively.The proposed approach achieves a good balance between accuracy and performance, comparing to the state-of-the-art methods. We attribute this to the Mamba’s capabilities in global modeling and autoregression, as well as its linear complexity, which offers greater potential for real-world deployment on resource-constrained devices, and provides valuable insights that may influence ongoing and future developments in this rapidly evolving field of optical flow.
Limitations.
While MambaFlow achieves results comparable to or superior to some state-of-the-art methods, there is still room for future improvement, as shown from the KITTI results in Table 1. Initially, it might not generalize well when there exists a substantial disparity between the training data and the test data. Secondly, it can still provide better accuracy under adverse conditions, such as motion blur, fog, and high-resolution, which remain unexplored. We will pretrain our model on large-scale datasets such as TartanAir [48] and VIPER [41], which can escalate its generalization ability. Furthermore, there are numerous outstanding variants of Mamba [63] have emerged recently, enabling us to adjust the entire architecture and its core Mamba-based components for more accurate estimations and to facilitate real-world applications.